# LLM4SBR: A Lightweight and Effective Framework for Integrating Large Language Models in Session-based Recommendation

### v1

@@ Abstract.
Traditional session-based recommendation (SBR) utilizes session behavior sequences from anonymous users for recommendation. Although this strategy is highly efficient, it sacrifices the inherent semantic information of the items, making it difficult for the model to understand the true intent of the session and resulting in a lack of interpretability in the recommended results. Recently, large language models (LLMs) have flourished across various domains, offering a glimpse of hope in addressing the aforementioned challenges. Inspired by the impact of LLMs, research exploring the integration of LLMs with the Recommender system (RS) has surged like mushrooms after rain. However, constrained by high time and space costs, as well as the brief and anonymous nature of session data, the first LLM recommendation framework suitable for industrial deployment has yet to emerge in the field of SBR.
To address the aforementioned challenges, we have proposed the LLM Integration Framework for SBR (LLM4SBR). Serving as a lightweight and plug-and-play framework, LLM4SBR adopts a two-step strategy. Firstly, we transform session data into a bimodal form of text and behavior. In the first step, leveraging the inferential capabilities of LLMs, we conduct inference on session text data from different perspectives and design the component for auxiliary enhancement. In the second step, the SBR model is trained on behavior data, aligning and averaging two modal session representations from different perspectives. Finally, we fuse session representations from different perspectives and modalities as the ultimate session representation for recommendation. We conducted experiments on two real-world datasets, and the results demonstrate that LLM4SBR significantly improves the performance of traditional SBR models and is highly lightweight and efficient, making it suitable for industrial deployment.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
The 21st century is an era of information explosion. Recently, Recommender systems (RS) (Lu et al., 2015; He et al., 2017) have received widespread attention from industry and academia as a crucial tool to solve information overload. To achieve personalized and accurate recommendations, RS usually uses the user’s personal information and historical behavior records to model user portraits. As both user privacy concerns and businesses’ demands for accurately capturing user dynamic intent continue to escalate, research on Session-based Recommendation (SBR) (Hidasi et al., 2015; Wu et al., 2019) has become a crucial aspect of RS, which only relies on the behavior sequence generated by the user within the session time to model, and does not require user profiles. Traditional SBR research (Adomavicius and Tuzhilin, 2005; Rendle et al., 2010; Hidasi et al., 2015; Wu et al., 2019; Wang et al., 2019) is based on the ID-based (behavior-based) recommendation paradigm. After encoding the items into one-hot features, methods such as Markov chains (Norris, 1998), recurrent neural networks (RNN) (Mandic and Chambers, 2001), graph neural networks (GNN) (Scarselli et al., 2008), etc., are employed to model anonymous session sequences.
While traditional recommendation methods can efficiently and accurately model collaborative information, they often overlook semantic information in interaction behavior, such as item name, price, etc. Particularly in SBR, where sequence lengths are typically short and data sparsity is high, and the SBR model uses one-hot encoding of the item ID to represent the item, which results in a serious lack of correlation between items. Consequently, solely modeling sparse behavioral information is insufficient for understanding users’ true intent. Semantic information, unlike interaction information, inherently possesses similarities and correlations between items. For example, if a user clicks on ”iPhone 15,” ”running shoes,” ”iPhone 14,” ”milk,” and ”skirt,” solely modeling behavior might mistakenly prioritize the last few clicks, such as ”milk” and ”skirt,” in determining session intent. However, leveraging semantic information, we can analyze that the user in the current session is likely more interested in the Apple product series. Therefore, if we can appropriately infer semantic information within the session, we can better understand the true intent behind the session sequence.
With the strong debut of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Zeng et al., 2022; Bai et al., 2023), it has not only shaken the entire field of natural language processing (NLP) (Chowdhary and Chowdhary, 2020), but also caused turmoil in various fields, and the RS is no exception. LLMs trained on large-scale corpora exhibit robust language comprehension abilities as well as a certain degree of logical reasoning capability. LLMs like GPT-4 are capable of handling complex tasks and engaging in dialogues, inspiring researchers in RS to envision new directions for the future development of RS. But how should RS be combined with LLM? This is the issue that RS researchers are most concerned about. Recently, there has been a proliferation of work exploring the combination of RS and LLM. Some works (Cui et al., 2022; Geng et al., 2022) take advantage of LLM by converting tasks in RS into language understanding or language generation tasks in NLP, through pre-training, fine-tuning, etc. Some researchers (Dai et al., 2023; Kang et al., 2023; Harte et al., 2023) have applied LLM to different recommendation system processes to explore LLM’s ability to encode features, sort, and score. The above attempts to combine RS with LLM have achieved many encouraging results, but they are not suitable for SBR. SBR combined with LLM has the following difficulties:
  • LLM hallucinations are more likely to occur. The sequence length in SBR is typically short, and access to users’ personal information is unavailable. When the information available to the LLM is very limited, the LLM may not be able to generate valid answers or may generate false items that exceed the item set returned.
  • A ”repeater” problem occurred when fine-tuning LLM. Session data is typically augmented through sequence splitting, resulting in datasets containing numerous similar sessions. Consequently, fine-tuning LLMs may lead to instances where the model excessively duplicates input text or repetitively repeats the same sentences when generating responses.
  • Training and inferring consume a lot of resources. LLM is complex in calculation takes up a lot of GPU memory, and takes a long time to infer. The recommendation task pursues real-time performance, so LLM-based RS models are difficult to implement in industrial practice.
@@
To address these difficulties, we propose a lightweight and effective LLM-enhanced framework framework (LLM4SBR) for SBR. The framework comprises two distinct stages, intent inference and representation enhancement. In the intent inference stage, our framework employs LLM as the inference engine. We guide LLM in inferring through carefully crafted prompts from different perspectives. The intent localization module is crafted to eliminate hallucinations and semantically enhance the reasoning results. Subsequently, these refined outcomes are encoded into an embedded form and stored in external files. Moving on to the representation enhancement stage, the traditional SBR model simultaneously loads interaction data in ID format and intention inference data in text format. On one hand, the SBR model models conversation representations from different perspectives based on interaction data, while on the other hand, it parses text data into embedded forms. Subsequently, alignment and uniformity of session embeddings and inference embeddings are performed separately for each perspective. Finally, all embeddings from all perspectives are fused as the ultimate session representation for prediction.
We summarize significant contributions as follows:
  • We are the first to propose an LLM enhancement framework for SBR. We divide the LLM inference and SBR model training into two stages. The LLM inference results are saved in an external file in advance, ensuring that GPU usage and training time during training depend only on the SBR model.
@@
  • We proposed an intent localization module, which addresses hallucinations and enhances semantic aspects in the preliminary results of LLM inference. In addition, We achieve a finer-grained modal alignment by performing alignment and uniformity between embeddings from different perspectives, facilitating the effective integration of interaction ID information and textual information.
  • Experiments on two real-world datasets show that our proposed framework LLM4SBR can be applied to most current SBR models and achieve substantial performance improvements.
@@ 2.1. Session-based Recommendation
In the field of SBR, the available information is very limited, consisting only of interaction data within the session. Therefore, the focus of SBR research lies in how to effectively model interaction behavior and learn session preferences. Based on different modeling emphases, we can broadly categorize SBR methods into two types: traditional SBR methods and methods focusing on modeling item transition relationships.
In traditional SBR methods, S-POP (Adomavicius and Tuzhilin, 2005) recommends based on the most popular items, and Item-KNN (Davidson et al., 2010) calculates item similarity based on historical behavior to recommend similar items. As Markov chains exhibit advantages in modeling sequential data, FPMC (Rendle et al., 2010) captures data sequence information and user preferences by combining first-order Markov chains with matrix factorization. In the SBR methods based on deep learning, inspired by the field of NLP, GRU4Rec (Hidasi et al., 2015) proposed for the first time to use of the recurrent neural network (RNN) to simulate user preference changes in behavioral sequence data. Based on this research, Stamp improved performance by introducing an attention mechanism to make preferences more targeted. NARM (Li et al., 2017) uses the attention network to capture users’ short-term interests and long-term dependencies.
As GNN shows its prowess in various fields, SBR researchers have found that by constructing session data into the form of graphs, they can better capture the complex transformation relationships between items and greatly improve recommendation performance. SR-GNN(Wu et al., 2019) is the first model to represent sequences in the form of session graphs, utilizing gated graph neural networks as encoders. GC-SAN (Xu et al., 2019), an upgraded version of SR-GNN, incorporates attention mechanisms to make session representations more targeted. GCE-GNN (Wang et al., 2020), HADCG (Su et al., 2023), MSGAT (Qiao et al., 2023) and KMVG(Chen et al., 2023) construct multiple graphs with different structures, simultaneously considering both local item collaborations and global session collaboration relationships. In addition, DHCN (Xia et al., 2021), HL(Wang et al., 2022a), and HIDE (Li et al., 2022) captures the complex high-order miscellaneous information of the items by building the hypergraph.
Although the aforementioned SBR methods have achieved good performance, they solely rely on modeling the interaction information of sessions, thus lacking effective utilization of semantic information embedded within the sequences.
@@ 2.2. Recommender System with LLM
Generative dialogue models represented by ChatGPT have caused a stir in research in various fields. According to how LLM participates in the recommendation system, we simply divide it into LLM as Recommender and LLM-enhanced Recommender.
@@ 2.2.1. LLM as Recommender
The model of LLM as Recommender realizes the transformation from the ID paradigm to the modal paradigm by converting the recommendation task into a task in natural language processing.
The M6-Rec (Cui et al., 2022) model extends the pre-trained language model M6 (Lin et al., 2021) by transforming recommendation tasks into either language understanding or language generation tasks. It establishes a unified foundational recommendation model to reduce downstream tasks’ dependence on data. Shijie Geng et al. (Geng et al., 2022) proposed the P5 paradigm, which enables predictions in a zero-shot or few-shot manner by providing adaptive personalized prompts tailored to different users. This approach reduces the need for fine-tuning. Wang-cheng Kang et al. (Kang et al., 2023) evaluated the performance of LLMs of different sizes (250M - 540B parameters) in zero-shot, few-shot, and fine-tuning scenarios to explore the extent to which LLM understands user preferences based on the user’s previous behavior. Sunhao Dai et al. (Dai et al., 2023) enhance the recommendation capabilities of ChatGPT by combining ChatGPT with traditional information retrieval (IR) ranking functions. GPT4Rec (Li et al., 2023d) first generates queries based on a language model, and then optimizes product retrieval separately through a search engine, addressing optimization from two aspects. VIP5 (Geng et al., 2023) explores a multi-modal base model of the P5 recommendation paradigm that considers both visual and textual modalities. Zhu Sun et al. (Sun et al., 2023) proposed the PO4ISR model of SBR, which promotes LLM to continuously reflect and update the results from the perspective of real-time optimization prompts to improve the accuracy of recommendations. Agent4Rec (Zhang et al., 2023) utilizes a generative agent empowered by LLM to simulate and infer personalized user preferences and behavioral patterns. The core of the above method is to enhance recommendation performance by improving LLM’s adaptability to recommended data and reasoning capabilities.
Although these methods have made breakthrough progress in zero-shot, few-shot, and interpretability aspects, they suffer from drawbacks such as high fine-tuning costs and difficulty in capturing specific fine-grained behavioral patterns. Consequently, they face challenges in being deployed in industrial applications.
@@ 2.2.2. LLM-enhanced Recommender
LLM-enhanced RS treats LLM as a tool to enhance the performance of recommendation models. The large-model recommendation framework proposed by Weiwei et al. (Wei et al., 2023) leverages graph-enhanced strategies based on LLM to enhance RS, addressing challenges posed by data sparsity and low-quality side information in RS. Chat-Rec (Gao et al., 2023) integrates traditional RS with conversational AI like ChatGPT, eliminating the need for training to gain a deep understanding of user preferences through LLM’s comprehension of dialogue context. CTRL (Li et al., 2023a) regards the original table data and the corresponding text data as two different modalities, and uses the collaborative CTR model and the pre-trained language model respectively for feature extraction, and adjusts the knowledge of the two modalities through comparative learning. LlamaRec retrieves candidates based on user interaction history through a sequence recommendation model. Candidates and historical records are designed as textual prompts, with the output of LLM transformed into a ranked probability distribution. E4SRec (Li et al., 2023b) is a solution that combines sequence recommendation with LLMs. It takes only ID sequences as input and ensures efficient controllable generation by predicting all candidate sequences at each forward pass. The above method has made us realize the potential of integrating LLM with RS and how a two-stage framework can better balance efficiency and performance compared to an end-to-end framework. Jesse Harte et al. (Harte et al., 2023) devised three strategies for leveraging LLM, and found that using embeddings initialized with LLM significantly enhances the performance of sequence recommendation models. This inspires us about the importance of textual semantics.
The above methods explore the effectiveness of LLM in RS from different perspectives. Compared to the ”LLM as Recommender” approach, they have greatly improved performance and efficiency. However, these methods do not fully leverage and integrate textual and interaction information, and they are not applicable to SBR scenarios with short sequences and no user information.
Considering aspects such as performance, efficiency, and hallucinations of LLMs, effectively integrating LLM with traditional RS models remains a challenging issue. In comparison, our framework is the first plug-and-play LLM framework designed for SBR. It effectively addresses several aspects mentioned above and better meets real-world industrial demands.

### v2

# Multi-view Intent Learning and Alignment with Large Language Models for Session-based Recommendation

@@ Abstract.
Session-based recommendation (SBR) methods often rely on user behavior data, which can struggle with the sparsity of session data, limiting performance. Researchers have identified that beyond behavioral signals, rich semantic information in item descriptions is crucial for capturing hidden user intent. While large language models (LLMs) offer new ways to leverage this semantic data, the challenges of session anonymity, short-sequence nature, and high LLM training costs have hindered the development of a lightweight, efficient LLM framework for SBR.
To address the above challenges, we propose an LLM-enhanced SBR framework that integrates semantic and behavioral signals from multiple views. This two-stage framework leverages the strengths of both LLMs and traditional SBR models while minimizing training costs. In the first stage, we use multi-view prompts to infer latent user intentions at the session semantic level, supported by an intent localization module to alleviate LLM hallucinations. In the second stage, we align and unify these semantic inferences with behavioral representations, effectively merging insights from both large and small models. Extensive experiments on two real datasets demonstrate that the LLM4SBR framework can effectively improve model performance. We release our codes along with the baselines at https://github.com/tsinghua-fib-lab/LLM4SBR.
@@
copyright: ccjournal: TOISjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3719344ccs: Information systems Recommender systems
@@ 1. Introduction
With the widespread application of Recommender systems (RS) (Lu et al., 2015; He et al., 2017; Yuan et al., 2024b, a) in multiple fields, traditional RS methods are no longer sufficient to meet users’ actual needs in real-world scenarios. Specifically, on the one hand, for the new-coming user, it is difficult to collect enough user behavior data and preference information to make accurate recommendations and encounter the ”cold start” problem. On the other hand, with the enhancement of awareness of data privacy protection globally, such as California’s California Consumer Privacy Act (CCPA) (Illman and Temple, 2019), which reinforces the protection of individual privacy rights, placing compliance pressure on traditional recommendation strategies that rely on user personal information. Therefore, against this backdrop, research on Session-based Recommendation (SBR) (Wu et al., 2019) technology is gradually becoming a hotspot, which aims to provide recommendations based on users’ transient anonymous session behavioral data with less risk to user privacy.
Traditional research in SBR relies heavily on deep neural networks to model user interaction sequences. This includes the use of Recurrent Neural Networks (RNNs) (Mandic and Chambers, 2001) with Attention Mechanisms to capture long-term dependencies between items in a session, as well as the use of Graph Neural Networks (GNNs) (Scarselli et al., 2008) and Hypergraph Neural Networks (HGNNs) (Feng et al., 2019) to model complex item transformation relationships in a session. Although these deep learning-based SBR models (Adomavicius and Tuzhilin, 2005; Rendle et al., 2010; Hidasi et al., 2015; Wu et al., 2019; Wang et al., 2019) show high effectiveness in modeling behavioral information, they are still faced with the challenge of high sparsity of session data. Specifically, the SBR model uses one-hot encoding to represent item IDs, greatly weakening the correlation between items. Therefore, it is difficult for the model to deeply analyze and grasp the user’s fundamental interests and potential intentions when only analyzing sparse behavior records. On the other hand, unlike interactive data, semantic data revolves around items’ inherent similarity and relevance, which may bring opportunities for the above challenge. For instance, assuming a user sequentially interacts with “iPhone 15, running shoes, iPhone 14, milk, skirt”, traditional SBR models may interpret recent interactions with “milk” and “skirt” as indicative of the user’s prevailing preferences. However, through a deeper semantic lens, it becomes apparent that the selection of “iPhone 15” and “iPhone 14” underscores a more enduring and profound interest in Apple-branded products. This underscores the potential of semantic information integration within RSs to more accurately discern and capture the underlying needs and users’ interests.
Currently, the rapid development of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Zeng et al., 2022; Bai et al., 2023) has made remarkable achievements in the depth and breadth of language understanding and shows a strong logical inference ability, opening up a new path for recommendation algorithms. Recent work (Cui et al., 2022; Geng et al., 2022) has begun to utilize pre-training and fine-tuning techniques to integrate core tasks in recommendation systems, transforming them into language understanding and generation problems in natural language processing (NLP). Through this approach, LLMs can generate personalized recommendation lists based on a comprehensive understanding of users’ past behaviors and personal preferences, providing detailed recommendations and significantly enhancing both user experience and recommendation accuracy. However, the high performance of LLM is often accompanied by high computational costs and latency issues, which largely limits their widespread application in real-time industrial scenarios. Therefore, ID-based collaborative information modeling methods still maintain their unique advantages and indispensable position. In addition, due to the high sparsity and anonymity of session data, research on incorporating LLM into SBR scenarios is still in its infancy. Current research mainly utilizes the capabilities of LLM by fine-tuning parameters(Bao et al., 2023a, b) or prompt optimization(Sun et al., 2023). The above work has conducted a preliminary exploration, with unresolved key challenges as follows:
  • Behavioral information based on the ID paradigm is the core of personalized recommender systems, which is even ignored by some existing methods solely based on semantic information.
  • Session data is often augmented using sequence segmentation, but this approach may result in a large number of similar session sequences in the generated dataset, which affects the fine-tuned LLMs based on this data to produce repetitive sentences when generating responses.
  • Fine-tuning and pre-training often require a large number of computational resources, which is difficult to deploy in real-world scenarios practically.
  • The high sparsity of the session data makes LLMs fail to generate valid answers or may generate incorrect items that are outside the candidate set.
@@
To address these challenges, we propose a lightweight and effective LLM-enhanced framework (LLM4SBR), which comprises two distinct stages: intent inference and representation enhancement. In the intent inference stage, we employ LLM as the inference engine by guiding LLM to infer through carefully designed prompts from different views of user intents. The intent localization module is crafted to alleviate hallucinations and semantically enhance the inference results. In the representation enhancement stage, we introduce the traditional SBR model to load interaction behaviors simultaneously, combined with textual intention inference results. On the one hand, traditional models generate conversation representations from different views based on interaction data, while on the other hand, it parses text data into embedded forms. Subsequently, each view performs alignment and uniformity of session and inference embeddings separately. Finally, we fuse all embeddings from all views as the session representation for prediction.
We summarize the contributions of this work as follows:
  • We take the pioneering step to propose a general framework for introducing large language models to the session-based recommendation, aiming to leverage LLM’s powerful logical-reasoning capability to facilitate deep semantic information integration. The framework ingeniously separates LLM’s inference process from the training of the traditional SBR model, forming a two-stage strategy, thereby achieving a balanced optimization of efficiency and effectiveness.
@@
  • We first propose an intent localization module, which can alleviate LLM hallucination and enhance semantic-level intent in preliminary results of LLM inference. We then achieve a finer-grained modal alignment by performing alignment and uniformity between embeddings from different views, facilitating the effective integration of interaction ID information and semantic information.
  • We conduct extensive experiments on two real-world datasets, and the results show that our proposed LLMSBR framework can be a plug-in solution that can steadily improve the performance of almost all the existing SBR models. Further studies verify the effectiveness of each component in our proposed LLMSBR framework.
@@ 2.1. Session-based Recommendation
The available information in the field of SBR is very limited, consisting only of interaction data within the session. Therefore, SBR research focuses on how to effectively model interaction behavior and learn session preferences. Based on different modeling emphases, we can broadly categorize SBR methods into two types: traditional SBR methods and methods focusing on modeling item transition relationships.
In traditional SBR methods, S-POP (Adomavicius and Tuzhilin, 2005) recommends based on the most popular items, and Item-KNN (Davidson et al., 2010) calculates item similarity based on historical behavior to recommend similar items. As Markov chains exhibit advantages in modeling sequential data, FPMC (Rendle et al., 2010) captures data sequence information and user preferences by combining first-order Markov chains with matrix factorization. In the SBR methods based on deep learning, inspired by the field of NLP, GRU4Rec (Hidasi et al., 2015) proposed for the first time to use of the RNN to simulate user preference changes in behavioral sequence data. Based on this research, Stamp improved performance by introducing an attention mechanism to make preferences more targeted. NARM (Li et al., 2017) uses the attention network to capture users’ short-term interests and long-term dependencies. DSAN(Yuan et al., 2021) adaptively filters the noisy information in the sequences through a two-layer sparse attention network, ensuring that the model can focus on items that truly reflect the user’s interests.
As GNNs show their prowess in various fields, SBR researchers have found that by constructing session data into the form of graphs, they can better capture the complex transformation relationships between items and greatly improve recommendation performance. SR-GNN(Wu et al., 2019) is the first model to represent sequences in the form of session graphs, utilizing gated graph neural networks as encoders. GC-SAN (Xu et al., 2019), an upgraded version of SR-GNN, incorporates attention mechanisms to make session representations more targeted. GCE-GNN (Wang et al., 2020), MSGAT (Qiao et al., 2023), MEGAN (Wang et al., 2023) and KMVG(Chen et al., 2023) construct multiple graphs with different structures, simultaneously considering both local item collaborations and global session collaboration relationships. In addition, DHCN (Xia et al., 2021), HL(Wang et al., 2022a), and HIDE (Li et al., 2022) captures the complex high-order miscellaneous information of the items by building the hypergraph. In recent research, GSNIR (Jin et al., 2023) adopts a dual intent network to learn user intent from the attention mechanism and historical data distribution respectively. MiaSRec (Choi et al., 2024) represents various session intents by deriving multiple session representations centered around each item and dynamically selecting important representations.
Although the aforementioned SBR methods have achieved good performance, they mainly focus on modeling the interaction information in the session but fail to fully explore and utilize the rich semantic information contained in the sequence.
@@ 2.2. Recommender System with LLM
Generative dialogue models represented by ChatGPT have sparked research in various fields. According to how LLM participates in the RS, we divide it into LLM as Recommender and LLM-enhanced Recommender.
@@ 2.2.1. LLM as Recommender
The model of LLM as Recommender realizes the transformation from the ID paradigm to the modal paradigm by converting the recommendation task into a task in natural language processing. The M6-Rec (Cui et al., 2022) model extends the pre-trained language model M6 (Lin et al., 2021) by transforming recommendation tasks into either language understanding or language generation tasks. It establishes a unified foundational recommendation model to reduce downstream tasks’ dependence on data. Geng et al. (Geng et al., 2022) proposed the P5 paradigm, which enables predictions in a zero-shot or few-shot manner by providing adaptive personalized prompts tailored to different users. This approach reduces the need for fine-tuning. Kang et al. (Kang et al., 2023) evaluated the performance of LLMs of different sizes (250M - 540B parameters) in zero-shot, few-shot, and fine-tuning scenarios to explore the extent to which LLM understands user preferences based on the user’s previous behavior. Sunhao Dai et al. (Dai et al., 2023) enhance the recommendation capabilities of ChatGPT by combining ChatGPT with traditional information retrieval (IR) ranking functions. GPT4Rec (Li et al., 2023d) first generates queries based on a language model, and then optimizes product retrieval separately through a search engine, addressing optimization from two aspects. VIP5 (Geng et al., 2023) explores a multi-modal base model of the P5 recommendation paradigm that considers both visual and textual modalities. Zhu Sun et al. (Sun et al., 2023) proposed the PO4ISR model of SBR, which promotes LLM to continuously reflect and update the results from the view of real-time optimization prompts to improve the accuracy of recommendations. Agent4Rec (Zhang et al., 2023) utilizes a generative agent empowered by LLM to simulate and infer personalized user preferences and behavioral patterns. ToolRec (Zhao et al., 2024) uses LLM as a proxy user to guide the recommendation process and call external tools to generate recommendation lists that are closely related to the user’s subtle preferences. Lin et al. (Lin et al., 2024) achieve the goal of efficiently fine-tuning LLM-based RS by introducing a data-pruning task.
Although these methods have made breakthrough progress in zero-shot, few-shot, and interpretability aspects, the core of the above method is to enhance recommendation performance by improving LLM’s adaptability to recommended data and inference capabilities. Therefore, they suffer from drawbacks such as high fine-tuning costs and difficulty capturing specific fine-grained behavioral patterns.
@@ 2.2.2. LLM-enhanced Recommender
LLM-enhanced RS treats LLM as a tool to enhance the performance of recommendation models. The large-model recommendation framework proposed by Weiwei et al. (Wei et al., 2024) leverages graph-enhanced strategies based on LLM to enhance RS, addressing challenges posed by data sparsity and low-quality side information in RS. Chat-Rec (Gao et al., 2023) integrates traditional RS with conversational AI like ChatGPT, eliminating the need for training to gain a deep understanding of user preferences through LLM’s comprehension of dialogue context. CTRL (Li et al., 2023a) regards the original table data and the corresponding text data as two different modalities, uses the collaborative CTR model and the pre-trained language model, respectively, for feature extraction, and adjusts the knowledge of the two modalities through comparative learning. E4SRec (Li et al., 2023b) is a solution that combines sequence recommendation with LLMs. It takes only ID sequences as input and ensures efficient controllable generation by predicting all candidate sequences at each forward pass. The above method has made us realize the potential of integrating LLM with RS and how a two-stage framework can better balance efficiency and performance compared to an end-to-end framework. Jesse Harte et al. (Harte et al., 2023) devised three strategies for leveraging LLM, and found that using embeddings initialized with LLM significantly enhances the performance of sequential recommendation models. This inspires us about the importance of textual semantics. The SAID (Hu et al., 2024) shows how to effectively use LLM to convert item IDs into semantically rich embedding vectors for downstream recommendation task models.
The above methods explore the effectiveness of LLM in RS from different views, allowing us to realize the potential of integrating LLM with RS and how the two-stage framework can better balance efficiency and performance compared with the end-to-end framework. However, these methods fail to fully integrate and fuse semantic information and behavioral data, and therefore perform poorly when dealing with SBR scenarios with high sparsity and lack of user information.
Table 1. Description of notations.
Notations Descriptions
$S$ The set of all session sequences
$s_{t}$ The session sequence of clicks at timestamp $t$
$I$ The set of all items
$i_{t,k}$ The $-k\textit{-}$th item clicked in the session at timestamp $t$
$Y$, $y$, $\hat{y}$ The label set of all sessions, the ground truth and predicted label of a single session
$\text{Text}_{\text{infer}}$, $\text{Text}_{\text{item}}$ LLM inference results and item names in text form
$E$, $e$ Text embeddings obtained using a pre-trained language model
$H$, $h$ The set of latent representations for all sessions and the latent representation for a single session
$d$,$W$,$b$ Dimension, weight matrix and biases
$T$ Alignment transformation matrix
$K$ The number of items that users are actually interested in
$r$ The number of items with similar semantics
$\tau$ Weight of alignment and uniformity loss functions
$\mathcal{L}$ Loss Function
3. Problem Formulation
The objective of SBR is to predict the next interaction item expected to occur in the current session history of an anonymous user. Here, we provide the problem definition in mathematical terms. Each data entry in the dataset represents a session sequence. Let the collection of all sessions be denoted as $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{m}\}$, where $m$ is the total number of sessions. The item set is the summary of items that have appeared in all sessions, which we define as =$\mathcal{I}=\{i_{1},i_{2},\cdots,i_{n}\}$, where $n$ is the total number of items in the set. We represent the t-th session $s_{t}$ in the dataset as $s_{t}=\{i_{t,1},i_{t,2},\cdots,i_{t,k},\cdots,i_{t,|s_{t}|}\}$, where $|s_{t}|$ is the length of the current session, and $i_{t,k}\in\mathcal{I}$ represents the $k$-th clicked item in the current session $s_{t}$. Based on the above symbols and descriptions, we define the modeling goal of session $s_{t}$ as predicting the click of the $|s_{t}|+1$th item based on the historical behavior records of $s_{t}$. The symbols and explanations involved in this article are listed in Table 1. In the following text, different symbols are distinguished by subscripts.